# Invariant Subspace Decomposition

### v1

@@ Appendix C Auxiliary results
Let $\mathcal{N}\subseteq\mathbb{N}$ and let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an orthogonal partition w.r.t. $(X_{t})_{t\in\mathcal{N}}$.333 This is defined analogously to an orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$. Then, there exists a joint block diagonalizer of $(\Sigma_{t})_{t\in\mathcal{N}}$. More precisely, there exists an orthonormal matrix $U\in\mathbb{R}^{p\times p}$ such that for all $t\in\mathcal{N}$ the matrix $\tilde{\Sigma}_{t}\coloneqq U^{\top}\Sigma_{t}U$ is block diagonal with $q$ diagonal blocks $\widetilde{\Sigma}_{t,j}=(U^{S_{j}})^{\top}\Sigma_{t}U^{S_{j}}$, $j\in\{1,\dots,q\}$ and of dimension $|S_{j}|=\operatorname{dim}(\mathcal{S}_{j})$, where $S_{j}\subseteq\{1,\dots,p\}$ indexes a subset of the columns of $U$. Moreover, $\Pi_{\mathcal{S}_{j}}=U^{S_{j}}(U^{S_{j}})^{\top}$.
@@
Let $U=(u_{1},\ldots,u_{p})\in\mathbb{R}^{p\times p}$ be an orthonormal matrix with columns $u_{1},\ldots,u_{p}$ such that for all $j\in\{1,\ldots,q\}$ there exists $S_{j}\subseteq\{1,\ldots,p\}$ such that $\mathcal{S}_{j}=\operatorname{span}(\{u_{k}\mid k\in S_{j}\})$. Such a matrix $U$ can be constructed by selecting an orthonormal basis for each of the disjoint subspaces $\{\mathcal{S}_{j}\}_{j=1}^{q}$. Furthermore, assume that the columns of $U$ are ordered in such a way that for all $i,j\in\{1,\ldots,q\}$ with $i<j$ it holds for all $k\in S_{i}$ and $\ell\in S_{j}$ that $k<\ell$. As the matrix $U$ is orthogonal, it holds for all $j\in\{1,\ldots,q\}$ that the projection matrix $\Pi_{\mathcal{S}_{j}}$ can be expressed as $\Pi_{\mathcal{S}_{j}}=U^{S_{j}}(U^{S_{j}})^{\top}$ and hence using the definition of orthogonal partition (see 6) it holds that, for all $t\in\mathcal{N}$,
@@
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$. Then it holds for all $t\in[n]$ that
$\gamma_{0,t}=\sum_{j=1}^{q}\operatorname{Var}\left(\Pi_{\mathcal{S}_{j}}X_{t}\right)^{\dagger}\operatorname{Cov}\left(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t}\right)$
and, for all $j\in\{1,\ldots,q\}$,
@@
Let $U\in\mathbb{R}^{p\times p}$ be an orthonormal matrix such that, for all $t\in[n]$, $\tilde{\Sigma}_{t}\coloneqq U^{\top}\Sigma_{t}U$ is block diagonal with $q$ diagonal blocks $\widetilde{\Sigma}_{t,1},\dots,\widetilde{\Sigma}_{t,q}$ of dimensions $\operatorname{dim}(\mathcal{S}_{1}),\dots,\operatorname{dim}(\mathcal{S}_{q})$. Such a matrix exists by Lemma 2 and each diagonal block is given by $\widetilde{\Sigma}_{t,j}\coloneqq(U^{S_{j}})^{\top}\Sigma_{t}U^{S_{j}}$ and the projection matrix $\Pi_{\mathcal{S}_{j}}$ can be expressed as $\Pi_{\mathcal{S}_{j}}=U^{S_{j}}(U^{S_{j}})^{\top}$. Now, for an arbitrary $t\in[n]$ it holds using the linear model (1) that $\gamma_{0,t}=\Sigma_{t}^{-1}\operatorname{Cov}(X_{t},Y_{t})$ and hence we use $U$ to get the following expansion
@@
In the last equality we have used that $\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})^{\dagger}=\Pi_{\mathcal{S}_{j}}\Sigma_{t}^{-1}\Pi_{\mathcal{S}_{j}}$, by Lemma 1. This completes the proof of Lemma 3. ∎
Lemma 4.
Let $\mathcal{N}\subseteq\mathbb{N}$ and let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an orthogonal partition w.r.t. $(X_{t})_{t\in\mathcal{N}}$. Then it holds for all $j\in\{1,\dots,q\}$ and for all $t\in\mathcal{N}$ that
$\operatorname*{arg\,max}_{\beta\in\mathcal{S}_{j}}\operatorname{\Delta Var}_{t}(\beta)=\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})^{\dagger}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t})=\Pi_{\mathcal{S}_{j}}\gamma_{0,t}.$ (28)
Moreover, if, in addition, $|\mathcal{N}|<\infty$, it holds for all $j\in\{1,\ldots,q\}$ that
$\operatorname*{arg\,max}_{\beta\in\mathcal{S}_{j}}\overline{\operatorname{\Delta Var}}(\beta)=\left(\frac{1}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)^{\dagger}\frac{1}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t}),$
where $\overline{\operatorname{\Delta Var}}(\beta)\coloneqq\frac{1}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{\Delta Var}_{t}(\beta)$.
@@
where $I_{p}$ denotes the identity matrix. Let $U\in\mathbb{R}^{p\times p}$ and $(\tilde{\Sigma}_{t})_{t\in\mathcal{N}}$ be defined for the orthogonal partition w.r.t. $(X_{t})_{t\in\mathcal{N}}$ as in Lemma 2. We now observe that
@@
For all $w\in\mathcal{S}_{j}$, this expression equals $\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})^{\dagger}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t})$. For all $w\not\in\mathcal{S}_{j}$, it is not in $\mathcal{S}_{j}$. It now suffices to show that $\Pi_{\mathcal{S}_{j}}\gamma_{0,t}=\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})^{\dagger}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t})$, which follows from the following computation
where the first and third equality follow from Lemma 3.
We now observe that for all $j\in\{1,\dots,q\}$ and for all $\beta\in\mathcal{S}_{j}$ it follows from (29) that
$\overline{\operatorname{\Delta Var}}(\beta)=\frac{2}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Cov}(Y_{t},\Pi_{\mathcal{S}_{j}}X_{t})\beta-\beta^{\top}\frac{1}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\beta$
which has gradient
$\nabla(\overline{\operatorname{\Delta Var}})(\beta)=\frac{2}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t})-\frac{2}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\beta.$
$\beta^{*}=\left(\frac{1}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)^{\dagger}\frac{1}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t}).$
By Lemma 1, it holds that $\left(\frac{1}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)^{\dagger}=\Pi_{\mathcal{S}_{j}}\left(\frac{1}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Var}(X_{t})\right)^{-1}\Pi_{\mathcal{S}_{j}}$ and therefore $\beta^{*}\in\mathcal{S}_{j}$. In order to apply Lemma 1, we in particular use that $U$ is such that, for all $t\in\mathcal{N}$, $\tilde{\Sigma}_{t}=U^{\top}\Sigma_{t}U$ is block diagonal with diagonal blocks given, for all $j\in\{1,\dots,q\}$, by $\tilde{\Sigma}_{t,j}=(U^{S_{j}})^{\top}\Sigma_{t}U^{S_{j}}$. This implies that $U^{\top}\left(\sum_{t\in\mathcal{N}}\Sigma_{t}\right)U$ is also block diagonal with diagonal blocks $(U^{S_{j}})^{\top}\left(\sum_{t\in\mathcal{N}}\Sigma_{t}\right)U^{S_{j}}$. Moreover, its inverse is block diagonal and, by the properties of orthogonal matrices, its diagonal blocks are $(U^{S_{j}})^{\top}\left(\sum_{t\in\mathcal{N}}\Sigma_{t}\right)^{-1}U^{S_{j}}$. It now remains to show that this is also the only solution in $\mathcal{S}_{j}$. All other solutions in $\mathbb{R}^{p}$ are given, for an arbitrary vector $w\in\mathbb{R}^{p}$, by
$\beta^{*}+\left(I_{p}-\left(\frac{1}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)^{\dagger}\left(\frac{1}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)\right)w$
$=$ $\beta^{*}+\left(I_{p}-\Pi_{\mathcal{S}_{j}}\right)w.$
The equality follows from the following computation
$\quad\left(\frac{1}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)^{\dagger}\left(\frac{1}{|\mathcal{N}|}\sum_{t\in\mathcal{N}}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)$
$=\Pi_{\mathcal{S}_{j}}\left(\sum_{t\in\mathcal{N}}\operatorname{Var}(X_{t})\right)^{-1}\Pi_{\mathcal{S}_{j}}\Pi_{\mathcal{S}_{j}}\left(\sum_{t\in\mathcal{N}}\operatorname{Var}(X_{t})\right)\Pi_{\mathcal{S}_{j}}$
$=U^{S_{j}}(U^{S_{j}})^{\top}\left(\sum_{t\in\mathcal{N}}\Sigma_{t}\right)^{-1}U^{S_{j}}(U^{S_{j}})^{\top}\left(\sum_{t\in\mathcal{N}}\Sigma_{t}\right)U^{S_{j}}(U^{S_{j}})^{\top}$
$=U^{S_{j}}\left(\sum_{t\in\mathcal{N}}\tilde{\Sigma}_{t}\right)^{-1}\left(\sum_{t\in\mathcal{N}}\tilde{\Sigma}_{t}\right)(U^{S_{j}})^{\top}$
$=\Pi_{\mathcal{S}_{j}},$
where $U$ is the matrix introduced in (31). The first equality follows again from Lemma 1. For all $w\in\mathcal{S}_{j}$, $\beta^{*}+(I_{p}-\Pi_{\mathcal{S}_{j}})w$ equals $\beta^{*}$. For all $w\not\in\mathcal{S}_{j}$, $\beta^{*}+(I_{p}-\Pi_{\mathcal{S}_{j}})w$ is not in $\mathcal{S}_{j}$. This concludes the proof for Lemma 4. ∎
@@
The first equality holds by Lemma 4, since $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ is indeed an orthogonal subspace partition w.r.t. $(X_{t})_{t\in[n]}$, see Lemma 6. The second equality follows from the definition of $\mathcal{S}^{\operatorname{inv}}$ and can be proved by Lemma 3 and, observing that the set of subspaces $\{\mathcal{S}_{j}\mid j\in\{1,\dots,q\}:\mathcal{S}_{j}\text{ opt-invariant on}[n]\}$ is an orthogonal partition w.r.t. $(\Pi_{\mathcal{S}^{\operatorname{inv}}}X_{t})_{t\in[n]}$. The third equality holds by Lemma 4 and the fourth by definition of an opt-invariant subspace on $[n]$. ∎
Lemma 6.
$\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ is an orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$.
Proof.
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be a fixed irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ according to which $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ are defined. By orthogonality of the subspaces in the partition and by definition of $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$, $\mathcal{S}^{\operatorname{res}}=(\mathcal{S}^{\operatorname{inv}})^{\perp}$. Moreover, by definition of orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$, it holds that
@@
The uniqueness of the spaces $\mathcal{U}_{j}$ then implies the uniqueness of the irreducible joint block diagonalizer $U$ up to block permutations and block-wise isometric transformations. To show this result, we now use Murota et al. (2010, Proposition 1). In particular, if the set $\{A_{t}\}_{t=1}^{n}$ contains at least one matrix $A$ with all distinct eigenvalues then the assumptions of Proposition $1$ are satisfied. Let $\{\mathcal{Q}_{1},\dots,\mathcal{Q}_{p}\}$ be the set of eigenspaces of $A$ and for all $i\in\{1,\dots,p\}$ let $m_{i}\coloneqq\operatorname{dim}(\mathcal{Q}_{i})=1$. The proposition then implies that for all $i\in\{1,\dots,p\}$ there exists $j\in\{1,\dots,q\}$ such that $\mathcal{Q}_{i}\subseteq\mathcal{U}_{j}$. Moreover, for all $i$ such that $\mathcal{Q}_{i}\subseteq\mathcal{U}_{j}$ it holds that $m_{i}=\bar{m}_{j}$. As a consequence, we obtain that, since the eigenvalues of $A$ are distinct, $m_{1}=\dots=m_{p}=1$ and therefore for all $j\in\{1,\dots,q\}$, $\bar{m}_{j}=1$. This concludes the proof for Lemma 7.
@@
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$. Then, for all $j\in\{1,\dots,q\}$ such that $\mathcal{S}_{j}$ is opt-invariant on $[n]$, it holds for all $t\in[n]$ that
@@
where we used the definition of opt-invariance on $[n]$ for the second equality and Lemma 4 for the third and fourth equality. Since the result holds for all $t\in[n]$, it also follows that
@@
The last equality follows from the definition of an orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$. ∎
@@
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ and let $\mathcal{S}^{\operatorname{inv}}$ be the corresponding invariant subspace. Moreover, let $\overline{\operatorname{Var}}(X)\coloneqq\frac{1}{n}\sum_{t=1}^{n}\operatorname{Var}(X_{t})$ and $\overline{\operatorname{Cov}}(X,Y)\coloneqq\frac{1}{n}\sum_{t=1}^{n}\operatorname{Cov}(X_{t},Y_{t})$. Finally, let $U^{\operatorname{inv}}$ be the submatrix of an arbitrary irreducible joint block diagonalizer $U$ corresponding to the irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ whose columns span $\mathcal{S}^{\operatorname{inv}}$. Then,
@@
Appendix D Proofs
D.1 Proof of Theorem 1

### v2

@@ Appendix D Auxiliary results
Let $\mathcal{N}\subseteq\mathbb{N}$ and let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an orthogonal and $(X_{t})_{t\in\mathcal{N}}$-decorrelating partition. Then, there exists a joint block diagonalizer of $(\Sigma_{t})_{t\in\mathcal{N}}$. More precisely, there exists an orthonormal matrix $U\in\mathbb{R}^{p\times p}$ such that for all $t\in\mathcal{N}$ the matrix $\tilde{\Sigma}_{t}\coloneqq U^{\top}\Sigma_{t}U$ is block diagonal with $q$ diagonal blocks $\widetilde{\Sigma}_{t,j}=(U^{S_{j}})^{\top}\Sigma_{t}U^{S_{j}}$, $j\in\{1,\dots,q\}$ and of dimension $|S_{j}|=\operatorname{dim}(\mathcal{S}_{j})$, where $S_{j}\subseteq\{1,\dots,p\}$ indexes a subset of the columns of $U$. Moreover, $\Pi_{\mathcal{S}_{j}}=U^{S_{j}}(U^{S_{j}})^{\top}$.
@@
Let $U=(u_{1},\ldots,u_{p})\in\mathbb{R}^{p\times p}$ be an orthonormal matrix with columns $u_{1},\ldots,u_{p}$ such that for all $j\in\{1,\ldots,q\}$ there exists $S_{j}\subseteq\{1,\ldots,p\}$ such that $\mathcal{S}_{j}=\operatorname{span}(\{u_{k}\mid k\in S_{j}\})$. Such a matrix $U$ can be constructed by selecting an orthonormal basis for each of the disjoint subspaces $\{\mathcal{S}_{j}\}_{j=1}^{q}$. Furthermore, assume that the columns of $U$ are ordered in such a way that for all $i,j\in\{1,\ldots,q\}$ with $i<j$ it holds for all $k\in S_{i}$ and $\ell\in S_{j}$ that $k<\ell$. As the matrix $U$ is orthogonal, it holds for all $j\in\{1,\ldots,q\}$ that the projection matrix $\Pi_{\mathcal{S}_{j}}$ can be expressed as $\Pi_{\mathcal{S}_{j}}=U^{S_{j}}(U^{S_{j}})^{\top}$ and hence using the definition of orthogonal partition (see (6)) it holds that, for all $t\in\mathcal{N}$,
@@
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition. Then it holds for all $t\in[n]$ and, for all $j\in\{1,\ldots,q\}$, that
@@
Let $U\in\mathbb{R}^{p\times p}$ be an orthonormal matrix such that, for all $t\in[n]$, $\tilde{\Sigma}_{t}\coloneqq U^{\top}\Sigma_{t}U$ is block diagonal with $q$ diagonal blocks $\widetilde{\Sigma}_{t,1},\dots,\widetilde{\Sigma}_{t,q}$ of dimensions $\operatorname{dim}(\mathcal{S}_{1}),\dots,\operatorname{dim}(\mathcal{S}_{q})$. Such a matrix exists by Lemma 4 and each diagonal block is given by $\widetilde{\Sigma}_{t,j}\coloneqq(U^{S_{j}})^{\top}\Sigma_{t}U^{S_{j}}$ and the projection matrix $\Pi_{\mathcal{S}_{j}}$ can be expressed as $\Pi_{\mathcal{S}_{j}}=U^{S_{j}}(U^{S_{j}})^{\top}$. The statement then follows from Lemma 3. ∎
Proof of Lemma 1
Proof.
Let $U\in\mathbb{R}^{p\times p}$ be an orthonormal matrix such that, for all $t\in[n]$, $\tilde{\Sigma}_{t}\coloneqq U^{\top}\Sigma_{t}U$ is block diagonal with $q$ diagonal blocks $\widetilde{\Sigma}_{t,1},\dots,\widetilde{\Sigma}_{t,q}$ of dimensions $\operatorname{dim}(\mathcal{S}_{1}),\dots,\operatorname{dim}(\mathcal{S}_{q})$. Such a matrix exists by Lemma 4 and each diagonal block is given by $\widetilde{\Sigma}_{t,j}\coloneqq(U^{S_{j}})^{\top}\Sigma_{t}U^{S_{j}}$ and the projection matrix $\Pi_{\mathcal{S}_{j}}$ can be expressed as $\Pi_{\mathcal{S}_{j}}=U^{S_{j}}(U^{S_{j}})^{\top}$.
Now, for an arbitrary $t\in[n]$ it holds using the linear model (1) that $\gamma_{0,t}=\Sigma_{t}^{-1}\operatorname{Cov}(X_{t},Y_{t})$ and hence we use $U$ to get the following expansion
@@
In the second to last equality, we have used that $\Pi_{\mathcal{S}_{j}}$ is a projection matrix and thus idempotent. In the last equality we have used that $\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})^{\dagger}=\Pi_{\mathcal{S}_{j}}\Sigma_{t}^{-1}\Pi_{\mathcal{S}_{j}}$, by Lemma 5. It now suffices to show that $\Pi_{\mathcal{S}_{j}}\gamma_{0,t}=\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})^{\dagger}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t})$, which follows from the following computation
where the first equality follows from the first part of this proof and and the third equality follows from Lemma 5. ∎
Proof of Lemma 2
@@
where $I_{p}$ denotes the identity matrix. Let $U\in\mathbb{R}^{p\times p}$ and $(\tilde{\Sigma}_{t})_{t\in\mathcal{N}}$ be defined for the orthogonal and $(X_{t})_{t\in\mathcal{N}}$-decorrelating partition as in Lemma 4. We now observe that
@@
For all $w\in\mathcal{S}_{j}$, this expression equals $\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})^{\dagger}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t})$. For all $w\not\in\mathcal{S}_{j}$, it is not in $\mathcal{S}_{j}$. This concludes the proof of Lemma 2. ∎
Lemma 6.
$\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ is an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition.
Proof.
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be a fixed irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition according to which $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ are defined. By orthogonality of the subspaces in the partition and by definition of $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$, $\mathcal{S}^{\operatorname{res}}=(\mathcal{S}^{\operatorname{inv}})^{\perp}$. Moreover, by definition of orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition, it holds that
@@
The first equality holds by Lemma 2, since $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ is indeed an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition, see Lemma 6. The second equality follows from the definition of $\mathcal{S}^{\operatorname{inv}}$ and can be proved by Lemma 1 and observing that the set of subspaces $\{\mathcal{S}_{j}\mid j\in\{1,\dots,q\}:\mathcal{S}_{j}\text{ opt-invariant on}[n]\}$ is an orthogonal and $(\Pi_{\mathcal{S}^{\operatorname{inv}}}X_{t})_{t\in[n]}$-decorrelating partition. The third equality holds by Lemma 2 and the fourth by definition of an opt-invariant subspace on $[n]$. ∎
@@
The uniqueness of the spaces $\mathcal{U}_{j}$ then implies the uniqueness of the irreducible joint block diagonalizer $U$ up to block permutations and block-wise isometric transformations. To show this result, we now use Murota et al. (2010, Proposition 1). In particular, if the set $\{A_{t}\}_{t=1}^{n}$ contains at least one matrix $A$ with all distinct eigenvalues then the assumptions of Proposition $1$ are satisfied. Let $\{\mathcal{Q}_{1},\dots,\mathcal{Q}_{p}\}$ be the set of eigenspaces of $A$ and for all $i\in\{1,\dots,p\}$ let $m_{i}\coloneqq\operatorname{dim}(\mathcal{Q}_{i})=1$. The proposition then implies that for all $i\in\{1,\dots,p\}$ there exists $j\in\{1,\dots,q\}$ such that $\mathcal{Q}_{i}\subseteq\mathcal{U}_{j}$. Moreover, for all $i$ such that $\mathcal{Q}_{i}\subseteq\mathcal{U}_{j}$ it holds that $m_{i}=\bar{m}_{j}$. As a consequence, we obtain that, since the eigenvalues of $A$ are distinct, $m_{1}=\dots=m_{p}=1$ and therefore for all $j\in\{1,\dots,q\}$, $\bar{m}_{j}=1$. This concludes the proof for Lemma 8. ∎
@@
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition. Then, for all $j\in\{1,\ldots,q\}$ it holds that
$\operatorname*{arg\,max}_{\beta\in\mathcal{S}_{j}}\overline{\operatorname{\Delta Var}}(\beta)=\left(\frac{1}{n}\sum_{t=1}^{n}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)^{\dagger}\frac{1}{n}\sum_{t=1}^{n}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t}).$
In addition, for all $j\in\{1,\dots,q\}$ such that $\mathcal{S}_{j}$ is opt-invariant on $[n]$, it holds for all $t\in[n]$ that
@@
Let $U\in\mathbb{R}^{p\times p}$ and $(\tilde{\Sigma}_{t})_{t\in[n]}$ be defined for the orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition as in Lemma 4. For all $j\in\{1,\dots,q\}$ and for all $\beta\in\mathcal{S}_{j}$ it follows from (30) that
$\overline{\operatorname{\Delta Var}}(\beta)=\frac{2}{n}\sum_{t=1}^{n}\operatorname{Cov}(Y_{t},\Pi_{\mathcal{S}_{j}}X_{t})\beta-\beta^{\top}\frac{1}{n}\sum_{t=1}^{n}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\beta$
which has gradient
$\nabla(\overline{\operatorname{\Delta Var}})(\beta)=\frac{2}{n}\sum_{t=1}^{n}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t})-\frac{2}{n}\sum_{t=1}^{n}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\beta.$
$\beta^{*}=\left(\frac{1}{n}\sum_{t=1}^{n}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)^{\dagger}\frac{1}{n}\sum_{t=1}^{n}\operatorname{Cov}(\Pi_{\mathcal{S}_{j}}X_{t},Y_{t}).$
By Lemma 3, it holds that $\left(\frac{1}{n}\sum_{t=1}^{n}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)^{\dagger}=\Pi_{\mathcal{S}_{j}}\left(\frac{1}{n}\sum_{t=1}^{n}\operatorname{Var}(X_{t})\right)^{-1}\Pi_{\mathcal{S}_{j}}$ and therefore $\beta^{*}\in\mathcal{S}_{j}$. In order to apply Lemma 3, we in particular use that $U$ is such that, for all $t\in[n]$, $\tilde{\Sigma}_{t}=U^{\top}\Sigma_{t}U$ is block diagonal with diagonal blocks given, for all $j\in\{1,\dots,q\}$, by $\tilde{\Sigma}_{t,j}=(U^{S_{j}})^{\top}\Sigma_{t}U^{S_{j}}$. This implies that $U^{\top}\left(\sum_{t=1}^{n}\Sigma_{t}\right)U$ is also block diagonal with diagonal blocks $(U^{S_{j}})^{\top}\left(\sum_{t=1}^{n}\Sigma_{t}\right)U^{S_{j}}$. Moreover, its inverse is block diagonal and, by the properties of orthogonal matrices, its diagonal blocks are $(U^{S_{j}})^{\top}\left(\sum_{t=1}^{n}\Sigma_{t}\right)^{-1}U^{S_{j}}$. It now remains to show that this is also the only solution in $\mathcal{S}_{j}$. All other solutions in $\mathbb{R}^{p}$ are given, for an arbitrary vector $w\in\mathbb{R}^{p}$, by
$\beta^{*}+\left(I_{p}-\left(\frac{1}{n}\sum_{t=1}^{n}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)^{\dagger}\left(\frac{1}{n}\sum_{t=1}^{n}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)\right)w$
$=$ $\beta^{*}+\left(I_{p}-\Pi_{\mathcal{S}_{j}}\right)w.$
The equality follows from the following computation
$\quad\left(\frac{1}{n}\sum_{t=1}^{n}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)^{\dagger}\left(\frac{1}{n}\sum_{t=1}^{n}\operatorname{Var}(\Pi_{\mathcal{S}_{j}}X_{t})\right)$
$=\Pi_{\mathcal{S}_{j}}\left(\sum_{t=1}^{n}\operatorname{Var}(X_{t})\right)^{-1}\Pi_{\mathcal{S}_{j}}\Pi_{\mathcal{S}_{j}}\left(\sum_{t=1}^{n}\operatorname{Var}(X_{t})\right)\Pi_{\mathcal{S}_{j}}$
$=U^{S_{j}}(U^{S_{j}})^{\top}\left(\sum_{t=1}^{n}\Sigma_{t}\right)^{-1}U^{S_{j}}(U^{S_{j}})^{\top}\left(\sum_{t=1}^{n}\Sigma_{t}\right)U^{S_{j}}(U^{S_{j}})^{\top}$
$=U^{S_{j}}\left(\sum_{t=1}^{n}\tilde{\Sigma}_{t}\right)^{-1}\left(\sum_{t=1}^{n}\tilde{\Sigma}_{t}\right)(U^{S_{j}})^{\top}$
$=\Pi_{\mathcal{S}_{j}}.$
The first equality follows again from Lemma 3. For all $w\in\mathcal{S}_{j}$, $\beta^{*}+(I_{p}-\Pi_{\mathcal{S}_{j}})w$ equals $\beta^{*}$. For all $w\not\in\mathcal{S}_{j}$, $\beta^{*}+(I_{p}-\Pi_{\mathcal{S}_{j}})w$ is not in $\mathcal{S}_{j}$.
@@
where we used the definition of opt-invariance on $[n]$ for the second equality, Lemma 2 for the third equality and Lemma 1 for the fourth equality. Since the result holds for all $t\in[n]$, it also follows that
@@
The last equality follows from the definition of an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition.
@@
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be an irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition and let $\mathcal{S}^{\operatorname{inv}}$ be the corresponding invariant subspace. Moreover, let $\overline{\operatorname{Var}}(X)\coloneqq\frac{1}{n}\sum_{t=1}^{n}\operatorname{Var}(X_{t})$ and $\overline{\operatorname{Cov}}(X,Y)\coloneqq\frac{1}{n}\sum_{t=1}^{n}\operatorname{Cov}(X_{t},Y_{t})$. Finally, let $U^{\operatorname{inv}}$ be the submatrix of an arbitrary irreducible joint block diagonalizer $U$ corresponding to the irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition whose columns span $\mathcal{S}^{\operatorname{inv}}$. Then,
@@
Appendix E Proofs
E.1 Proof of Theorem 1